OpenStudy (anonymous):

9 HATBOX= 4 BOXHAT if these variables are being multiplied together, what is HATBOX/BOXHAT? explanation please

OpenStudy (anonymous):

...

hero (hero):

4/9

OpenStudy (anonymous):

explanation?

hero (hero):

replace hatbox = x boxhat = y 9y = 4x x/y = 4/9

hero (hero):

hatbox/boxhat = 4/9

hero (hero):

I don't know what else you could possibly put

OpenStudy (anonymous):

yes i understand that, but i don't really get how dividing the two gives you the answer.

hero (hero):

Trust me, that's the answer. If not you can embarrass me on here tomorrow when you find out what the right answer is

OpenStudy (anonymous):

deal.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i believe I've proven you wrong, unfortunately.

hero (hero):

How?

OpenStudy (anonymous):

well, hatbox and boxhat are the same variables, only in a different order, right?

hero (hero):

Oh, looks like you tricked me

hero (hero):

If that's the case, then no solution

OpenStudy (anonymous):

therefore, 9x=4y is not a true statement

hero (hero):

\[1 \ne \frac{4}{9}\]

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

so basically, i have no idea.

hero (hero):